OpenStudy (anonymous):

can someone help me find an equation for the inverse of the function. f(x)=2x -3^x/4

OpenStudy (anonymous):

Is this your equation? \( \huge f(x) = -3^{\frac{x}{4}} + 2x\) Or is it this one \( \huge f(x) = -\frac{3^{x}}{4} + 2x\)

OpenStudy (anonymous):

Or is this your equation \( \huge f(x) = \frac{-3^{x} + 2x}{4}\)